OpenStudy (anonymous):

The sum of the squares of three consecutive positive integers is 2030. What is the middle integer?

OpenStudy (anonymous):

(n)^2+(n+1)^2+(n+2)^2=2030

OpenStudy (unklerhaukus):

26

OpenStudy (anonymous):

And how did you come up with 26?

OpenStudy (anonymous):

n^2+n^2+2n+1+n^2+4n+4=2030 3n^2+6n+5=2030 3n^2+6n-2025=0 n^2+2n-675=0 (n-25)(n+27)=0 n=25, n=-27 Since it is positive the first number is 25. The second once is equal to 26 :)

OpenStudy (unklerhaukus):

well i used a guess and check method, i started with \[24^2+25^2+26^2=1877\] and then tried \[25^2+26^2+27^2=2030\] which turned out right.